the scale on a set of plans states that 1/8 inch equals 1 foot. if a water line measures 5 1/2 inches on the plans, what would its actual length be?
Answer by stanbon(75887)

the scale on a set of plans states that 1/8 inch equals 1 foot. if a water line measures 5 1/2 inches on the plans, what would its actual length be?
-------------------------
Use a proportion:
x/[5 1/2] = 1/[1/8]
x/(11/2) = 8
x = (11/2)*8
x = 44 feet
==================
Cheers,
Stan H.